Below are five test items, each belonging in one of the four cognitive levels of CAPS: knowledge, routine procedure, complex procedure or problem solving. Indicate in each case to what cognitive level the question belongs:

(a) What is the temperature difference between 17°C maximum – 3°C minimum?

(b) Ben walks to school in 48 minutes. Nomsa goes by bicycle in half the time that Ben walks. Koki goes by taxi in half the time that Nomsa goes by bicycle. At what time do they all start at home if they all arrive at school at 07:15?

(c)  How many centimeters are in a meter?

(d)  What is the mode of the data set below:

5; 6; 1; 4; 7; 3; 8; 2; 4; 3; 8; 9; 4; 6; 1; 2; 6; 5

(e) Calculate the area of the darker shaded outer part below. The larger rectangle is 15cm long and 9cm wide; the sides of the smaller square are 5cm each.

 

 



1
Expert's answer
2022-05-18T10:03:51-0400

(a)


"\Delta T=17\degree C-(-3\degree C)=20\degree C"


(b)

Ben walks to school in 48 minutes.

Nomsa goes by bicycle to school in 24 minutes.

Koki goes by taxi to school in 12 minutes.


Ben starts at home at 06:27.

Nomsa starts at home at 06:51.

Koki starts at home at 07:03.


(c) There are 100 centimeters in a meter


(d)


"1,1,2,2,3,3,4,4,4,5, 5, 6,6,6,7,8,8,9"

Mode "6, 4," each appeared 3 times.


(e)


"Area=15(9)-(5)^2=110({cm}^2)"